# Engineered Simultaneity: The Physical Impossibility of Consolidated Price Discovery Across Spacelike-Separated Exchanges

### 03 Mar 2026 v2

@@ Abstract
We define engineered simultaneity: the construction of a system that requires temporal comparison of events at spacelike-separated locations, implements this comparison via an implicit simultaneity convention, and represents the result as an objective measurement rather than a conventional choice. We show that the National Best Bid and Offer (NBBO)—the regulatory cornerstone of U.S. equity markets—is an instance of engineered simultaneity. The NBBO requires determining “current” prices across exchanges whose spatial separation places their price events outside each other’s light cones. Special relativity proves that the temporal ordering of such events is frame-dependent: there exist inertial reference frames in which the NBBO differs from the value reported by the Securities Information Processor. The impossibility is not approximate; it is exact and unavoidable within the causal structure of Minkowski spacetime. We formalize the argument using the causal precedence relation, connect it to Lamport’s theorem on distributed ordering, and note that approximately \$5 billion per year in latency arbitrage profits are extracted from the gap between the NBBO’s implicit simultaneity convention and physical reality.
@@ 6  Discussion
Three structural observations follow:
@@ 7  Conclusion
The NBBO assumes a spacelike hypersurface of simultaneity that special relativity proves cannot exist. We have formalized this as Theorem 1, connected it to Lamport’s impossibility of distributed ordering, and shown that the gap between the NBBO’s implicit simultaneity convention and physical reality is exploited for approximately \$5 billion annually.

### 23 Mar 2026 v3

@@ Abstract
We define engineered simultaneity: the construction of a system that requires temporal comparison of events at spacelike-separated locations, implements this comparison via an implicit simultaneity convention, and represents the result as an objective measurement rather than a conventional choice. We show that the National Best Bid and Offer (NBBO)—the regulatory cornerstone of U.S. equity markets—is an instance of engineered simultaneity. The NBBO requires determining “current” prices across exchanges whose spatial separation places their price events outside each other’s light cones. Special relativity proves that the temporal ordering of such events is frame-dependent: there exist inertial reference frames in which the NBBO differs from the value reported by the Securities Information Processor. The impossibility is not approximate; it is exact and unavoidable within the causal structure of Minkowski spacetime. General relativity compounds the impossibility: gravitational time dilation introduces frame-rate discrepancies between exchanges at different altitudes, and recent work on indefinite causal order in quantum information theory undermines the premise of a fixed causal structure altogether. We formalize the special-relativistic argument using the causal precedence relation, connect it to Lamport’s theorem on distributed ordering, and note that approximately \$5 billion per year in latency arbitrage profits are extracted from the gap between the NBBO’s implicit simultaneity convention and physical reality.
@@ 6  Discussion
Lewis [19] popularized the idea that markets were “rigged” by speed differentials and could be fixed by leveling the temporal playing field. The implicit assumption was that there exists a meaningful notion of a simultaneous market state that everyone could, in principle, reach. That assumption is incorrect. Flash Boys correctly identified latency arbitrage but misdiagnosed its root cause. The problem is not insufficient synchronization. The problem is the belief that synchronization can create simultaneity in the first place.
Four structural observations follow:
@@
Fourth, the impossibility extends beyond special relativity. General relativity further undermines the premise: clocks at different gravitational potentials tick at different rates—an effect actively corrected in GPS [24]. HFT infrastructure spans buildings, elevations, thermal gradients, and mechanical environments. Nanosecond agreement is already a negotiated approximation layered atop unprovable assumptions about symmetry and stability [22]. The idea of a single, authoritative market clock is physically indefensible.
6.1  Implications for Time Infrastructure
The analysis presented here was delivered as an invited talk at the Open Compute Project Time Appliances Project (OCP-TAP) on January 28, 2026 [27]. OCP-TAP builds precision time infrastructure for data centers and distributed systems. The relevance is direct: time appliances measure, but they do not define reality. Synchronization has limits that must be made explicit. Infrastructure should expose uncertainty, not hide it [28]. A companion paper [30] traces the fiction of synchronized time from its cultural origins through daylight saving time and leap seconds to its consequences in distributed computing and financial markets.
6.2  Beyond Relativity: Indefinite Causal Order
The impossibility argument can be strengthened further. In quantum mechanics, time is not an observable—there is no time operator. Time is a parameter external to the system, inferred from interaction rather than observed as a conserved quantity [26]. Protocols that treat timestamps as ground truth rather than metadata are therefore already on shaky physical footing.
Modern quantum information theory goes further. Work by Oreshkov, Costa, and Brukner [25] shows that some processes have no definite causal order at all. Event $A$ is not before or after event $B$ in any global sense. This is not philosophy; it is experimentally realizable physics. As markets become faster, more reflexive, and more tightly coupled, their behavior increasingly resembles systems where causal order is local and observer-dependent rather than globally defined.
6.3  The Structure of Exploitation
A system is functionally fraudulent when it claims physical legitimacy it cannot have, exploits asymmetries invisible to most participants, and converts uncertainty into private profit. Engineered simultaneity satisfies all three conditions. The HFT industry understands that simultaneity is not physically real, yet markets the appearance of it as neutral, objective, and fair. The real advantage is not speed—speed differences are visible and competitive. The real advantage is control over the definition of “now”: control over where time is measured, when clocks are sampled, and whose clock counts. This is not about speed. It is about time-definition control, and most market participants cannot audit it.
@@ 7  Conclusion
The NBBO assumes a spacelike hypersurface of simultaneity that special relativity proves cannot exist. General relativity compounds the problem through gravitational time dilation, and quantum information theory further undermines the premise through indefinite causal order. We have formalized the special-relativistic impossibility as Theorem 1, connected it to Lamport’s impossibility of distributed ordering, and shown that the gap between the NBBO’s implicit simultaneity convention and physical reality is exploited for approximately \$5 billion annually.
@@
A companion paper develops the broader theoretical framework—the Forward-In-Time-Only (FITO) category mistake in distributed systems [29]—while a second traces why synchronized time is itself a fiction, from its cultural origins through daylight saving time and leap seconds to its consequences in financial markets [30].